OpenStudy (experimentx):

Solve: \[ x^2 \left ( y -x {dy \over dx}\right ) = y \left ( dy\over dx\right)^2\] Soln: \[ y^2 = cx^2 + c^2 \]

OpenStudy (turingtest):

*bookmark!

OpenStudy (anonymous):

You're doing DEs now? Okay... I'll help you tomorrow if no one answers your question. But I'm tired rite now.

OpenStudy (anonymous):

multiply both sides of equation by \(y\) u will get \[ x^2y^2-x^3y \frac{dy}{dx}= (y \frac{dy}{dx})^2 \] suppose that \(t\) is a function of \(x\) let \[ y \frac{dy}{dx}=tx \] plug in the original equation : \(x^2y^2-tx^4= t^2x^2 \) or \(y^2=tx^2+ t^2 \) now take derivative of both sides we have \[2 y \frac{dy}{dx}=2xt=2xt+x^2 \frac{dt}{dx}+2t\frac{dt}{dx} \] or \[x^2 \frac{dt}{dx}+2t\frac{dt}{dx}=0 \]now we have 2 options ----> \(t'=0\) or \(2t=-x^2\) case \((I)\) ......... \(t'=0\) \(t=c\) or \(y\frac{dy}{dx}=cx\) and \(y^2=cx^2+d\) put back in this original equation we will get \(d=c^2\) so for this case answer is \(y^2=cx^2+c^2\) case \((II)\) ..........\(2t=-x^2\) \(2xt=-x^3\) or \(2y\frac{dy}{dx}=-x^3\) and from this case we get \[ y^2=-\frac{x^4}{4}+b\]substituting this in the original equation gives \(b=0\) so \(y^2=-x^4/4\) and this is not acceptable...

OpenStudy (anonymous):

***put back this in.............lol

OpenStudy (experimentx):

this is a guessing game!! right substitution would yield in to Clairaut's equation which isn't difficult to solve afterwards.

OpenStudy (experimentx):

the problems is how to think like the people who devise these kinds of problem ... i could never do this.

OpenStudy (anonymous):

me too...maybe i could never do this...i used the answer to get the right sub

OpenStudy (experimentx):

x^2 = u, y^2 = v for this

OpenStudy (experimentx):

i have another problem i'm having trouble with (have answer but no solution) .... i'll post in a while.